OpenStudy (anonymous):

An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth. 20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (sleepyjess):

\(V=\dfrac{4}{3}\pi r^2\)

OpenStudy (sleepyjess):

d=3.5 r=d/2 r=3.5/2

OpenStudy (sleepyjess):

does that help?

OpenStudy (anonymous):

yes, thank you so much!! :)

OpenStudy (sleepyjess):

for any of these volume questions you can always just google volume of a ...

OpenStudy (anonymous):

very true, thank you so much I really appreciate it

OpenStudy (sleepyjess):

\(\Huge\frak\color{skyblue}{No~Problem!}\)